OpenStudy (anonymous):

through (3,5) m=-3.. with medal

OpenStudy (kc_kennylau):

Point slope form first, coz you're given point and slope, then convert to whatever form you like :)

OpenStudy (haseeb96):

(x,y)=(3,5) m=-3 y=mx+b 5=-3(3)+b 5+9=b b=14 y=-3x+14

OpenStudy (kc_kennylau):

Point-slope form: \(y-5=-3(x-3)\) (using the information given) \[y-5=-3(x-3)\]\[y=-3x+9+5\]\(y=-3x+14\) (Slope-intercept form) \[y=-3x+14\]\(3x+y=14\) (Standard form)